OpenStudy (anonymous):

A baseball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 10. (h = height, t = time in seconds) Find the maximum height of the ball. a) 10 feet b) 26 feet c) 23 feet d) 2.3 feet

OpenStudy (anonymous):

Not sure if this is an optimization problem in your calculus class, so I'll answer from both an algebraic perspective and a calculus perspective. Perspective 1: Algebra your function gives the equation of a parabola of the form \[ax^2+bx+c\] where a=-16, b=32, and c=10. The crucial thing to recognize is that all parabolas are symmetric about the vertex. This means that if you find the x-intercepts, all you have to do is locate the point that is equidistant between them. this vertex will give you the maximum height. There is a simple formula for this (which is derived from the quadratic equation). \[x_{vertex}=-b/(2a)=-32/(2*-16)=1\] plugging this x value back into the original function gives the maximum height of \[-16+32+10=42-16=26\] Thus, the maximum height of the ball is 26 ft. Perspective 2: Calculus This is the exact same thing as perspective 1 except that you use differentiation to find the vertex formula. The derivation is below. set the equation of a parabola equal to zero \[ax^2+bx+c=0\] take the derivative \[2*ax+b=0\] solve for x \[x=-b/(2*a)\] And that's that :)

OpenStudy (anonymous):

thank you